OpenStudy (anonymous):

write the expression as a single logarithm (parentheses after log is log base) 4log(a) Y-1/3log(a) Z+5log(a) W

OpenStudy (campbell_st):

is the expression 1. \[\frac{4\log_{a}(y) - 1}{3\log_{a}(z)+5\log_{a}(w)}\] or 2. \[4\log_{a}(y) - \frac{1}{3} \log_{a}(z) + 5\log_{a}(w)\]

OpenStudy (anonymous):

expression 2

OpenStudy (campbell_st):

ok...you need to know about powers and logs \[\log(a^b) = b \times \log(a)\] so you need to rewrite each log with a power e.g. \[4 \log(y) = \log(y^4)\] the next one the power is 1/3 and the last log, the power is 5

OpenStudy (campbell_st):

next you need the laws for dividing and multiplying multiplying \[\log(ab) = \log(a) + \log(b)\] dividing \[\log(\frac{a}{b}) = \log(a) - \log(b)\]

OpenStudy (anonymous):

what I came up with is \[\log_{a} \frac{ y ^{4}w ^{5} }{ \sqrt[3]{z}}\] is this correct? am I missing the base of the log?

OpenStudy (campbell_st):

yep... that looks really good... well done the only thing, to avoid confusion it should be in brackets \[\log_{a} (...)\]

OpenStudy (anonymous):

brackets or parentheses?

OpenStudy (anonymous):

is there another way to write cube root at the bottom of a fraction?

OpenStudy (campbell_st):

brackets or parentheses.... doesn't matter... I'd use brackets... the only alternative for the denominator is index form..

OpenStudy (anonymous):

could I write \[Z ^{\frac{1 }{ 3}}\]

OpenStudy (campbell_st):

yep... thats fine... but for me... using the cube root... is neater..

OpenStudy (anonymous):

cool yeah I agree, but the online form im using doesnt have the option of writing it that way

OpenStudy (anonymous):

thanks for your help!

OpenStudy (campbell_st):

good luck